OpenStudy (anonymous):

A box contains 20 lightbulbs, of which 5 are defective. If 4 lightbulbs are picked from the box randomly, the probability that at most 2 of them are defective is ________?

OpenStudy (anonymous):

?

OpenStudy (kropot72):

The probabilities of 0, 1 and 2 defectives need to be calculated and added. \[\large P(0)=\frac{15C4}{20C4}\] \[\large P(1)=\frac{5C1\times 15C3}{20C4}\] \[\large P(2)=\frac{5C2\times 15C2}{20C4}\]

OpenStudy (anonymous):

okay thanks :) i think i got it from here

OpenStudy (kropot72):

Let me know if your need more help on this.